# Phase field modelling of the detachment of bubbles from a solid substrate

### v1

@@ 3 Transient evolution and stability/instability
[…]
@@
Figure 3: Numerical simulations of the evolution of a bubble with initial contact angle $53\pi/$ 5 and Bond number 4 (time in dimensionless units), below criticality.
@@
Figure 4: Numerical simulations of the evolution of a bubble with initial contact angle $53\pi/$ 5 and Bond number 15 (time in dimensionless units).
@@ 4 Inclined substrates
Angle $B_{o}$ (No slope) $B_{o}$ (inclined plane ${}^{\circ}$) $B_{o}$ (inclined plane ${}^{\circ}$)
@@ 5 The effect of external flows
In figure 9 we plot various stages of the evolution of the bubble under the action of an external velocity field as computed by the phase field method. The Bond number is 15 and the initial contact angle is $53\pi/$ 5. The inlet velocity is a laminar flow in the right yx-plane, given by $v=-Uz$, with $U=3$.
@@
Figure 9: Numerical simulations of the evolution of a bubble with initial contact angle $53\pi/$ 5, Bond number 10 (time in dimensionless units). There is an inlet linear flow $v_{y}=-Uz$ with $U=3$ in dimensionless units.

### v2

@@ 3 Transient evolution and stability/instability
$\int_{0}^{\delta h}\pi\frac{2y}{\kappa}dy=\frac{\pi}{\kappa}(\delta h)^{2}=\frac{1}{2}A_{GS}\delta h.$
@@
Figure 3: Numerical simulations of the evolution of a bubble with initial contact angle $3\pi/5$ and Bond number 4 (time in dimensionless units), below criticality.
@@
Figure 4: Numerical simulations of the evolution of a bubble with initial contact angle $3\pi/5$ and Bond number 15 (time in dimensionless units).
@@ 4 Inclined substrates
Angle $B_{o}$ (No slope) $B_{o}$ (inclined plane 30) $B_{o}$ (inclined plane 60)
@@ 5 The effect of external flows
In figure 9 we plot various stages of the evolution of the bubble under the action of an external velocity field as computed by the phase field method. The Bond number is 15 and the initial contact angle is $3\pi/5$. The inlet velocity is a laminar flow in the right yx-plane, given by v=Uzv=-Uz, with $U=3$.
@@
Figure 9: Numerical simulations of the evolution of a bubble with initial contact angle $3\pi/5$, Bond number 10 (time in dimensionless units). There is an inlet linear flow $v_{y}=-Uz$ with $U=3$ in dimensionless units.